OpenStudy (anonymous):

Use the Function to answer the question: f(x)=-2|x-5|+3 What is the vertex of the graph of f(x)? a. (-5,3) b. (-2,3) c.(5,-3) d. (5,3)

OpenStudy (jdoe0001):

notice the absolute value expression what value of "x" makes that absolute expression = 0?

OpenStudy (anonymous):

Hmmm well -5 is in the absolute value. would that be it?

OpenStudy (jdoe0001):

x = -5?

OpenStudy (jdoe0001):

well... let's say we give "x" some value what value given to "x" would make the absolute value expression = 0, I meant

OpenStudy (anonymous):

I'm horrible at math. but to my understanding... in order to find the value of "x" I need to see what I would plug in for it to see if that equals 0?

OpenStudy (anonymous):

In that case... I can honestly say I am not sure. I am sorry.

OpenStudy (jdoe0001):

let us look at the absolute value expression \(\large | x- 5|\) let us use say x = 7 so if we set x = 7 then | (7) -5 | => | 2 | => 2 so... setting x = 7 doesn't make the expression 0, it actually makes is 2 any ideas on what we could use for "x" to make the expression 0?

OpenStudy (anonymous):

keep loosing connection. Let me read what you said

OpenStudy (anonymous):

OH! Sorry. I understand what you're saying now. we could put 5 as the value of "x" to get 0 because 5-5=0

OpenStudy (jdoe0001):

yeap thus \(\large { f(x)=-2|x-5|+3 \begin{cases} -2|{\color{brown}{ 5}}-5|{\color{blue}{ +3}} \\ \quad \\ ({\color{brown}{ 5}},{\color{blue}{ +3}})\impliedby vertex \end{cases} }\)

OpenStudy (anonymous):

Oh, awesome!! Thanks so much jdoe001!!

OpenStudy (jdoe0001):

yw